OpenStudy (anonymous):

Can someone help me understand what I did wrong? What is the solution set for |2x – 3| = 13? 2x-3=13 +3 +3 2x=16 16/2=8 x=8 2x+3=13 -3 -3 2x=10 10/2=5 x=5

jimthompson5910 (jim_thompson5910):

it should be 2x-3 = 13 or 2x - 3 = -13

jimthompson5910 (jim_thompson5910):

not 2x+3 = 13

OpenStudy (campbell_st):

the 1st solution is correct... the 2nd solution has an incorrect sign its 2x +3 = -13

OpenStudy (anonymous):

Ooooh!! I see it now, thanks a looooot!!! :DD So now the answer is 8 or -8! Thanks a lot you two! c: xD

jimthompson5910 (jim_thompson5910):

no, one of those answers you got just now is wrong

OpenStudy (campbell_st):

no its 8 and 2x - 3 = -13 add 3 2x = -10

OpenStudy (anonymous):

Dang it, it's -5 or 8 then.. Dx

jimthompson5910 (jim_thompson5910):

yep, -5 or 8 is correct

OpenStudy (anonymous):

Finally, thank you very much. c:

jimthompson5910 (jim_thompson5910):

I recommend checking any/all answers you get

jimthompson5910 (jim_thompson5910):

so if you checked x = -8 (one wrong solution, but we'll check it anyway), then you would have gotten |2x – 3| = 13 |2*(-8) – 3| = 13 ... plug in x = -8 |-16 – 3| = 13 |-19| = 13 19 = 13 ... which is FALSE so x = -8 is NOT a solution

OpenStudy (anonymous):

Now you're just confusing me, like... It's the answer isn't it?

jimthompson5910 (jim_thompson5910):

I'm showing you a way to check for potential answers

jimthompson5910 (jim_thompson5910):

if I claim x = -8 is an answer, you can use this check method to disprove my claim

OpenStudy (anonymous):

;~; Oh, ok... Thank you.

jimthompson5910 (jim_thompson5910):

np